If X is a binomial random variable with expected value 6 and variance 2.4. Find the probability of X=5.
1
Expert's answer
2020-04-07T16:25:02-0400

"np=6" and "npq=2.4"

Thus,"q=\frac{2.4}{6}=0.4"

"p=1-q=1-0.4=0.6"

"n=\frac{6}{0.6}=10"

Thus, "P(X=5)=\binom{10}{5}×0.6^5×0.4^5"

"=252\u00d70.07776\u00d70.01024=0.20066"

Similarly, the Excel formula =BINOM.DIST(5,10,0.6,FALSE) can be used to obtain the answer.